# Parameter-Efficient Fine-Tuning in Spectral Domain for Point Cloud Learning

### v1

@@ 5.1 Datasets and Implement Details
TABLE II: The comparison between other fine-tuning strategies and our PointGST. We report the overall accuracy (OA) and trainable parameters (Params.) across three variants of ScanObjectNN [68] and ModelNet40 [79]. All methods use default data augmentation as the baseline. indicates reproduced results. ScanObjectNN results are reported without voting, while ModelNet40 results are presented both with and without voting (-/-).
@@
Point-PEFT [64] AAAI 24 0.7 (3.13%) - - - 85.00(+1.93) 93.4(+0.7) /    -
@@
5.2 Compared Methods
To illustrate the effectiveness of our approach, popular pre-train models like Point-BERT [85], Point-MAE [47], ACT [11], ReCon [50], and PointGPT [4] are selected as our baselines. Additionally, we include the SOTA point cloud PEFT methods for comparisons: 1) IDPT [88] utilizes DGCNN [74] to generate instance-aware prompts before the last transformer layer for model fine-tuning rather than relying on static prompts; 2) Point-PEFT [64] utilize a point-prior bank for feature aggregation and incorporating Adapters into each transformer block. 3) DA [15] uses a dynamic aggregation method to replace previous static aggregation for pre-trained point cloud Transformers. 4) DAPT [99], the latest SOTA point cloud PEFT method, proposes to adjust tokens based on significance scores.
It is important to note that among the methods mentioned above, IDPT and DAPT conduct more comprehensive experiments compared to the other two approaches. As a result, we primarily follow the experimental setups of IDPT and DAPT to organize our experiments.
6 Results and Analysis
6.1 3D Classification
6.1.1 The comparisons of different fine-tuning strategies
We first comprehensively compare the representative fine-tuning strategies used in the point cloud tasks, including the fully fine-tuning (FFT), IDPT [88], and DAPT [99], in terms of trainable parameters (Params.) and performance. We apply our PointGST to various pre-trained models, such as Point-BERT [85], Point-MAE [47], ACT [11], ReCon [50], and PointGPT-L [4], size ranging from 22.1M to 360.5M. Tab. II lists the detailed results, and it can find that:
TABLE III: Classification on ScanObjectNN PB_T50_RS [68] with strong data augmentation. Overall accuracy (%) without voting is reported. Params. is the trainable parameters.
@@ 6.1.1 The comparisons of different fine-tuning strategies
TABLE IV: Comparison of overall accuracy (OA) and trainable parameters (Params.) across three variants of ScanObjectNN [68] and ModelNet40 [79]. Results highlighted in green indicate the use of the voting strategy.
@@
1) PointGST effectively balances the number of trainable parameters and performance. Based on the quantitative comparisons, it is evident that increasing the scale of a pre-trained model will improve the performance. However, this would meanwhile come at a huge cost of fine-tuning, e.g., Point-MAE and PointGPT-L report 85.18% and 93.4% with 22.1M and 360.5M parameters on the ScanObjectNN PB_T50_RS dataset, respectively. Thanks to only fine-tuning a few parameters in the spectral domain, our method can effectively address this dilemma. Specifically, on the most convincing pre-trained point cloud models, PointGPT-L, compared with the fully fine-tuning (FFT), we only need 2.4M trainable parameters, which is merely 0.6% of the FFT. More importantly, with so few training parameters, we even outperform the FFT by 1.77%, 0.99%, 1.43%, and 0.7% on the OBJ_BG, OBJ_OBLY, PB_T50_RS and ModelNet40 datasets, respectively.
Compared to spatial domain-based methods like IDPT and DAPT, our PointGST achieves superior performance with fewer trainable parameters by leveraging the spectral domain, where features are more easily de-correlated. In addition, our PointGST introduces intrinsic information via the Graph Fourier Transform (GFT), enabling targeted fine-tuning. In short, PointGST proves to be a potential candidate that brings promising performance and parameter-efficient approaches to low-resource scenarios.
@@
2) PointGST can effectively generalize to diverse pre-trained models. A good PEFT method should consistently perform well across different pre-trained models, regardless of the pre-training strategies, data used, or model sizes. However, our observations reveal that existing PEFT methods [88, 99, 64, 15] fail to deliver consistent improvements across diverse pre-trained models. For instance, as shown in Tab. II, we apply five distinct pre-trained models, varying in pre-training techniques (e.g., masked modeling and contrastive learning) and sizes (ranging from 22.1M to 360.5M parameters). Notably, both IDPT and DAPT exhibit negative impacts when applied to baselines like ACT and ReCon. Even in scenarios where these methods produce positive outcomes, the gains from IDPT and DAPT remain limited. In contrast, our PointGST consistently yields improvements across most conducted pre-trained models and outperforms IDPT and DAPT.
In addition, we find that both Point-PEFT [64] and DA [15] apply the point cloud rotation as the data augmentation [11] on the Point-MAE. To keep fair, we also adopt the augmentations to conduct experiments. The comparisons in Tab. III clearly show that PointGST consistently achieves higher accuracy than Point-PEFT and DA.
6.1.2 Compared with state-of-the-art methods
The comparisons between the state-of-the-art (SOTA) methods and our proposed method on ScanObjectNN and ModelNet40 are illustrated in Tab. IV. We categorize the comparison methods into two types: supervised learning only and self-supervised representation learning with full fine-tuning. The results yield two notable observations regarding the efficacy of the proposed PointGST:
1) Self-supervised representation learning methods [4, 50, 11] generally outperform those based solely on supervised learning [55, 22, 44], emphasizing the critical role of pre-training. Additionally, these self-supervised methods often involve more trainable parameters than their supervised counterparts (e.g., 13.2M and 657.2M for PointMLP [44] and ReCon++-L [51], respectively), which underscores the practical relevance of parameter-efficient fine-tuning.
2) When using PointGPT-L [4] as the baseline, our PointGST outperforms all previous methods, establishing a new state-of-the-art, while requiring only 2.4M trainable parameters. To the best of our knowledge, this is the first approach to nearly saturate the performance on the ScanObjectNN OBJ_BG dataset (e.g., 99.48% OA). Note that it is not something that can be brought about by simply fitting a dataset. Instead, it highlights the current limitations of existing point cloud analysis datasets in effectively evaluating new methods. Consequently, we encourage the community to develop more challenging datasets to better assess the progress in future point cloud analysis research.
6.2 Few-shot Learning
Few-shot learning plays a crucial role in evaluating the efficiency of data usage for training. We make few-shot comparisons on ModelNet40 and the results are listed in Tab. V. Our PointGST beats the previous PEFT methods [88, 99] in all conducted settings. To further verify the few-shot training ability of our method, we conducted experiments on the challenging ScanObjectNN PB_T50_RS dataset, where we set the range from 2% to 100%, as depicted in Fig. 5. Specifically, our PointGST demonstrates strong robustness when trained on limited data. Especially with only 2% of the training data, PointGST achieves an overall accuracy of 81.96%, significantly outperforming IDPT [88] and DAPT [99] by 1.74% and 4.51%, respectively. In addition, we report the least trainable parameters, only 0.6M. The results indicate that PointGST efficiently leverages limited training data to capture the underlying characteristics of point clouds, outperforming existing point cloud PEFT methods, particularly in data-constrained environments.
Figure 6: The t-SNE visualizations, along with the trainable parameters (Params.) and overall accuracy (OA), are derived from the ScanObjectNN PB_T50_RS dataset using a pre-trained PointGPT-L with various tuning strategies. The final classification features for t-SNE visualizations are extracted from the last linear layer.
Figure 7: (a) Default point tokens $\boldsymbol{T}_{s}$. (b) Initializing a learnable matrix to act on $\boldsymbol{T}_{s}$, still in the spatial domain. (c) The proposed method transfers the $\boldsymbol{T}_{s}$ (graph signal) to the spectral tokens.
@@ 6.2 Few-shot Learning
TABLE V: Few-shot learning on ModelNet40 [79]. Overall accuracy±$\pm$±standard deviation without voting is reported.
@@
with Self-Supervised Representation Learning (Full fine-tuning)
MaskPoint [40] ECCV 22 95.0$\pm$3.7 97.2$\pm$1.7 91.4$\pm$4.0 93.4$\pm$3.5
Point-M2AE [93] NeurIPS 22 96.8$\pm$1.8 98.3$\pm$1.4 92.3$\pm$4.5 95.0$\pm$3.0
ACT [11] ICLR 23 96.8$\pm$2.3 98.0$\pm$1.4 93.3$\pm$4.0 95.6$\pm$2.8
PointGPT-L [4] NeurIPS 23 98.0$\pm$1.9 99.0$\pm$1.0 94.1$\pm$3.3 96.1$\pm$2.8
ReCon++-L [51] ECCV 24 98.0$\pm$2.3 99.5$\pm$0.8 94.5$\pm$4.1 96.5$\pm$3.0
with Self-Supervised Representation Learning (Efficient fine-tuning)
@@
6.3 Part Segmentation
Part segmentation presents challenges in accurately predicting detailed class labels for each point. Following DAPT [99], we apply the Point-BERT, Point-MAE, and ReCon as baselines on the ShapeNetPart dataset. We incorporated the proposed point cloud spectral adapter (PCSA) into every layer of the models. The quantitative results are shown in Tab. VI. Our method achieves highly competitive performance while utilizing significantly fewer trainable parameters than fully fine-tuning counterparts (i.e., baseline methods). Notably, the increase in trainable parameters is mainly due to the head part, but we still reduce overall trainable parameters and surpass the performance of other PEFT methods like IDPT [88], Point-PEFT [64] and DAPT [99].
TABLE VI: Part segmentation on the ShapeNetPart [83]. The mIoU for all classes (Cls.) and for all instances (Inst.) are reported. Params. represents the trainable parameters. denotes reproduced results.
@@ 6.3 Part Segmentation
6.4 Scene-level Point Cloud Analysis
We evaluate the proposed PointGST on the semantic segmentation task using the S3DIS dataset [2], with results shown in Tab. VII. Our empirical observations suggest that previous PEFT methods [88, 64, 99] struggle with this task, likely due to the models being pre-trained on object-level datasets [3]. This results in sub-optimal performance when fine-tuned on a scene-level dataset. In contrast, the proposed PointGST effectively utilizes intrinsic information during fine-tuning, bridging the performance gap between fully and efficiently fine-tuning strategies. For example, using ACT [11] as the baseline, PointGST achieves 67.6% mAcc and 57.4% mIoU with only 5.55M trainable parameters, substantially outperforming IDPT [88], Point-PEFT [64] and DAPT [99] by 3.5%/5.3%, 1.6%/2.8%, and 3.5%/2.9% in mAcc/mIoU, respectively. Similarly, when applied to Point-MAE [47] and ReCon [50], PointGST consistently surpasses both IDPT, Point-PEFT, and DAPT, further demonstrating its robustness across different pre-trained models.
TABLE VII: Semantic segmentation on the S3DIS [2]. The mean accuracy (mAcc) and mean IoU (mIoU) are reported. Params. represents the trainable parameters.

### v2

@@ 4.3 Insight of PEFT in the Spectral Domain
TABLE I: The comparison between other fine-tuning strategies and our PointGST. We report the overall accuracy (OA) and trainable parameters (Params.) across three variants of ScanObjectNN [75] and ModelNet40 [76]. All methods use default data augmentation as the baseline. ScanObjectNN results are reported without voting, while ModelNet40 results are presented both with and without voting (-/-).
@@
Point-PEFT [18] AAAI 24 0.7 (3.13%) 7.61 88.81(+1.38) 89.67(+1.55) 85.00(+1.93) 93.4(+0.7) /    -
@@
TABLE II: Classification on ScanObjectNN PB_T50_RS [75] with strong data augmentation. Overall accuracy (%) without voting is reported. Params. is the trainable parameters.
@@
5 EXPERIMENTAL Setup
5.1 Datasets and Implementation Details
All the experiments are conducted on a single 24 GB NVIDIA GPU. During fine-tuning, the parameters of the pre-trained backbones remain frozen, while only the newly added parameters are updated. The hyper-parameters, scaling factor $s$, group number $k$, and adapter dimension $r$ are set to 1, 4, and 36, respectively. For all conducted baselines, these hyper-parameters are consistently applied across all datasets/tasks without specific adjustments. We employ a transpose format of Z space-filling curve [77] to scan the key points to construct the local graphs.
ScanObjectNN [75] is a challenging real-world 3D object dataset with $\sim$15K indoor point cloud instances across 15 categories, including three variants (OBJ_BG, OBJ_ONLY, PB_T50_RS) of increasing difficulty. We follow baseline training settings: AdamW optimizer [78] (weight decay 0.05), cosine learning rate schedule [79] (initial LR 5e-4, 10-epoch warm-up), 300 epochs, batch size 32, and 2048 input points (divided into 128 groups of 32 points each).
ModelNet40 [76] is a synthetic dataset of 12,311 CAD models from 40 categories. The training follows ScanObjectNN settings, with 1024 input points (divided into 64 groups of 32 points each) and 300 epochs.
ShapeNetPart [80] is a point-level part segmentation benchmark with 16,881 instances from 16 objects and 50 part categories. We use AdamW (weight decay 1e-4), cosine scheduler (initial LR 1e-4, 10-epoch warm-up), 300 epochs, batch size 16, and 2048 points (divided into 128 groups of 32 points each).
S3DIS [81] contains six large-scale indoor areas with 273M points over 13 categories. Following [82], Area 5 is used for evaluation. Models are trained for 60 epochs (batch size 32, initial LR 2e-4, cosine scheduler). Other settings follow ShapeNetPart.
ScanNetV2 [83] comprises 1,201 training, 312 validation, and 100 test scenes with 18 annotated categories. We use AdamW (weight decay 0.1), cosine scheduler (initial LR 2e-4, 10-epoch warm-up), 1080 epochs, and batch size 8.
PCN [84] contains 30,974 pairs of point clouds from 8 categories. The network takes 2048 points as inputs for each and completes the other 14,336 points. We use AdamW (weight decay 5e-4), cosine scheduler (initial LR 1e-4), 100 epochs, and batch size 16.
5.2 Baseline Methods and Comparisons
To illustrate the effectiveness of our approach, popular pre-trained models like Point-BERT [9], Point-MAE [8], ACT [11], ReCon [37], and PointGPT [13] are selected as our baselines on all datasets. Additionally, we include the SOTA point cloud PEFT methods for comparisons: 1) IDPT [16] utilizes DGCNN [61] to generate instance-aware prompts before the last transformer layer for model fine-tuning rather than relying on static prompts; 2) Point-PEFT [18] utilizes a point-prior bank for feature aggregation and incorporating Adapters into each transformer block. 3) DA [19] uses a dynamic aggregation method to replace previous static aggregation for pre-trained point cloud Transformers. 4) DAPT [17], the latest SOTA point cloud PEFT method, proposes to adjust tokens based on significance scores. It should be noted that for datasets or tasks where certain methods have not provided results, we reproduce their performance using the official code to ensure a comprehensive comparison.
6 Results and Analysis
6.1 3D Classification
6.1.1 The comparisons of different fine-tuning strategies
We first comprehensively compare the representative fine-tuning strategies used in the point cloud tasks, including the fully fine-tuning (FFT), IDPT [16], Point-PEFT [18], and DAPT [17], in terms of trainable parameters (Params.) and performance. We apply our PointGST to various pre-trained models [9, 8, 11, 37, 13], size ranging from 22.1M to 360.5M. Tab. I lists the detailed results, and we found that:
TABLE III: Comparison of overall accuracy (OA) and trainable parameters (Params.) across three variants of ScanObjectNN [75] and ModelNet40 [76]. Results highlighted in green indicate the use of the voting strategy.
@@ 6.1.1 The comparisons of different fine-tuning strategies
1) PointGST effectively balances the number of trainable parameters and performance. Based on the quantitative comparisons, it is evident that increasing the scale of a pre-trained model will improve the performance. However, this would meanwhile come at a huge cost of fine-tuning, e.g., Point-MAE and PointGPT-L report 85.18% and 93.4% with 22.1M and 360.5M parameters on the ScanObjectNN PB_T50_RS dataset, respectively. Thanks to only fine-tuning a few parameters in the spectral domain, our method can effectively address this dilemma. Specifically, on the most convincing pre-trained point cloud models, PointGPT-L [13], compared with the fully fine-tuning (FFT), we only need 2.4M trainable parameters, which is merely 0.67% of the FFT. More importantly, with so few training parameters, we even outperform the FFT by 1.77%, 0.99%, 1.43%, and 0.7% on the OBJ_BG, OBJ_ONLY, PB_T50_RS and ModelNet40 datasets, respectively. Compared to spatial domain-based methods like IDPT, Point-PEFT, and DAPT, our PointGST achieves superior performance with fewer trainable parameters by leveraging the spectral domain, where features are more easily de-correlated. In short, PointGST proves to be a potential candidate that brings promising performance and parameter-efficient approaches to low-resource scenarios.
@@
2) PointGST can effectively generalize to diverse pre-trained models. A good PEFT method should consistently perform well across different pre-trained models, regardless of the pre-training strategies, data used, or model sizes. However, our observations reveal that existing PEFT methods [16, 17, 18, 19] fail to deliver consistent improvements across diverse pre-trained models. For instance, as shown in Tab. I, we apply five distinct pre-trained models, varying in pre-training techniques (e.g., masked modeling and contrastive learning) and sizes (ranging from 22.1M to 360.5M parameters). Notably, IDPT, Point-PEFT, and DAPT exhibit negative impacts when applied to baselines like ACT [11] and ReCon [37]. Even in scenarios where these methods produce positive outcomes, the gains from IDPT, Point-PEFT, and DAPT remain limited. In contrast, our PointGST consistently yields improvements across most conducted pre-trained models and outperforms existing point cloud PEFT methods. In addition, we find that both Point-PEFT [18] and DA [19] apply the point cloud rotation as the data augmentation [11] on the Point-MAE. To keep fair, we also adopt the augmentations to conduct experiments. The comparisons in Tab. II clearly show that PointGST consistently achieves higher accuracy than Point-PEFT and DA.
@@
TABLE IV: Few-shot learning on ModelNet40 [76]. Overall accuracy±$\pm$±standard deviation without voting is reported.
@@
ACT  [11] (baseline) ICLR 23 96.8$\pm$2.3 98.0$\pm$1.4 93.3$\pm$4.0 95.6$\pm$2.8
+ IDPT [16] ICCV 23 96.7$\pm$2.6 98.2$\pm$0.9 92.4$\pm$4.5 95.5$\pm$3.0
+ Point-PEFT [18] AAAI 24 95.4$\pm$2.7 97.7$\pm$1.7 91.8$\pm$5.1 94.9$\pm$3.0
+ DAPT [17] CVPR 24 94.4$\pm$3.9 97.3$\pm$2.4 90.6$\pm$6.1 95.1$\pm$3.7
+ PointGST (ours) - 97.2$\pm$1.9 98.6$\pm$1.5 92.8$\pm$4.0 95.5$\pm$3.1
ReCon  [37] (baseline) ICML 23 97.3$\pm$1.9 98.9$\pm$1.2 93.3$\pm$3.9 95.8$\pm$3.0
+ IDPT [16] ICCV 23 96.8$\pm$2.2 98.6$\pm$0.8 92.7$\pm$3.8 95.9$\pm$3.2
+ Point-PEFT [18] AAAI 24 95.4$\pm$2.6 97.7$\pm$1.4 91.5$\pm$5.0 95.2$\pm$3.4
+ DAPT [17] CVPR 24 95.6$\pm$3.4 97.3$\pm$2.0 91.9$\pm$4.9 94.5$\pm$3.2
+ PointGST (ours) - 96.9$\pm$2.2 98.7$\pm$0.9 92.9$\pm$3.8 95.8$\pm$2.8
PointGPT-L  [13] (baseline) NeurIPS 23 98.0$\pm$1.9 99.0$\pm$1.0 94.1$\pm$3.3 96.1$\pm$2.8
+ IDPT [16] ICCV 23 96.8$\pm$2.0 98.3$\pm$1.4 95.3$\pm$2.5 96.5$\pm$2.5
+ Point- PEFT [18] AAAI 24 95.7$\pm$2.5 98.0$\pm$1.8 94.8$\pm$2.7 96.3$\pm$2.4
+ DAPT [17] CVPR 24 97.2$\pm$3.4 98.4$\pm$1.3 95.9$\pm$2.7 97.3$\pm$2.4
+ PointGST (ours) - 97.4$\pm$2.0 99.6$\pm$0.5 96.3$\pm$3.6 97.6$\pm$2.9
6.1.2 Compared with state-of-the-art methods
The comparisons between the state-of-the-art (SOTA) methods and our proposed method on ScanObjectNN and ModelNet40 are illustrated in Tab. III. We categorize the comparison methods into two types: supervised learning only and self-supervised representation learning with full fine-tuning. The results yield two notable observations regarding the efficacy of the proposed PointGST:
1) Self-supervised representation learning methods [13, 37, 11] generally outperform those based solely on supervised learning [87, 89, 88], emphasizing the critical role of pre-training. Additionally, these self-supervised methods often involve more trainable parameters than their supervised counterparts (e.g., 13.2M and 657.2M for PointMLP [88] and ReCon++-L [15], respectively), which demonstrates the practical relevance of parameter-efficient fine-tuning.
2) When using PointGPT-L [13] as the baseline, our PointGST outperforms all previous methods, establishing a new state-of-the-art, while requiring only 2.4M trainable parameters. To the best of our knowledge, this is the first approach to nearly saturate the performance on the ScanObjectNN OBJ_BG dataset (e.g., 99.48% OA). Note that this is not attributable to simply overfitting the dataset. Instead, it highlights the current limitations of existing point cloud analysis datasets in effectively evaluating new methods. Consequently, we encourage the community to develop more challenging datasets to better assess the progress in future point cloud analysis research.
TABLE V: Part segmentation on the ShapeNetPart [80]. The mIoU for all classes (Cls.) and for all instances (Inst.) are reported. Params. represents the trainable parameters.
@@ 6.1.2 Compared with state-of-the-art methods
ACT [11] (baseline) ICLR 23 27.06 84.66 86.1
+ IDPT [16] ICCV 23 5.69 83.82 85.9
+ Point-PEFT [18] AAAI 24 5.62 81.25 84.4
+ DAPT [17] CVPR 24 5.65 83.44 85.5
+ PointGST (ours) - 5.59 84.04 85.8
@@
PointGPT-L [13] (baseline) NeurIPS 23 339.39 84.80 86.6
+ IDPT [16] ICCV 23 33.17 82.91 85.1
+ Point-PEFT [18] AAAI 24 31.94 82.34 84.9
+ DAPT [17] CVPR 24 33.85 83.03 85.2
+ PointGST (ours) - 31.85 83.91 85.8
6.2 Few-shot Learning
Few-shot learning plays a crucial role in evaluating the efficiency of data usage for training. We make few-shot comparisons on ModelNet40 and the results are listed in Tab. IV. Our PointGST beats the previous PEFT methods [16, 18, 17] in all conducted settings. To further verify the few-shot training ability of our method, we conducted experiments on the challenging ScanObjectNN PB_T50_RS dataset, where we set the range from 2% to 100%, as depicted in Fig. 5. Specifically, our PointGST demonstrates strong robustness when trained on limited data. Especially with only 2% of the training data, PointGST achieves an overall accuracy of 81.96%, significantly outperforming IDPT [16], Point-PEFT [18], and DAPT [17] by 1.74%, 4.34%, and 4.51%, respectively. In addition, we report the least trainable parameters, only 0.6M. The results indicate that PointGST efficiently leverages limited training data to capture the underlying characteristics of point clouds, outperforming existing point cloud PEFT methods, particularly in data-constrained environments.
TABLE VI: Semantic segmentation on the S3DIS [81]. The mean accuracy (mAcc) and mean IoU (mIoU) are reported. Params. represents the trainable parameters.
@@ 6.2 Few-shot Learning
Point-BERT [9] (baseline) CVPR 22 27.02 70.3 60.8
+ Linear probing - 5.20 64.1 53.9
+ IDPT [16] ICCV 23 5.64 65.2 54.2
+ Point-PEFT [18] AAAI 24 5.58 66.9 56.0
+ DAPT [17] CVPR 24 5.61 67.0 56.3
+ PointGST (ours) - 5.55 68.6 58.4
@@
PointGPT-L [13] (baseline) NeurIPS 23 339.30 70.6 62.2
+ Linear probing - 29.93 62.8 53.0
+ IDPT [16] ICCV 23 33.09 69.1 57.6
+ Point-PEFT [18] AAAI 24 31.86 68.4 58.0
+ DAPT [17] CVPR 24 33.76 68.7 59.0
+ PointGST (ours) - 31.76 69.1 59.6
TABLE VII: Scene-level object detection on the ScanNetV2 [83]. The average precision with 3D IoU thresholds of 0.25 ($AP_{25}$) and 0.5 ($AP_{50}$) are reported. Params. represents the trainable parameters.
Methods Reference Params. (M) $AP_{25}$ (%) $AP_{50}$ (%)
Point-BERT [9] (baseline) CVPR 22 34.63 61.0 38.3
+ Linear probing - 13.35 56.9 35.5
+ IDPT [16] ICCV 23 14.68 58.3 36.5
+ Point-PEFT [18] AAAI 24 13.81 58.1 36.2
+ DAPT [17] CVPR 24 13.99 58.3 37.0
+ PointGST (ours) - 13.71 58.4 38.6
Point-MAE [8] (baseline) ECCV 22 34.63 59.3 39.8
+ Linear probing - 13.35 58.9 38.1
+ IDPT [16] ICCV 23 14.68 58.9 38.2
+ Point-PEFT [18] AAAI 24 13.81 58.0 38.9
+ DAPT [17] CVPR 24 13.99 59.6 38.5
+ PointGST (ours) - 13.71 59.9 39.2
ACT [11] (baseline) ICLR 23 34.63 59.9 39.8
+ Linear probing - 13.35 56.2 36.2
+ IDPT [16] ICCV 23 14.68 58.1 36.2
+ Point-PEFT [18] AAAI 24 13.81 57.5 36.0
+ DAPT [17] CVPR 24 13.99 57.6 37.9
+ PointGST (ours) - 13.71 60.3 40.3
ReCon [37] (baseline) ICML 23 34.63 59.6 40.0
+ Linear probing - 13.35 57.6 37.6
+ IDPT [16] ICCV 23 14.68 57.4 36.2
+ Point-PEFT [18] AAAI 24 13.81 58.5 38.2
+ DAPT [17] CVPR 24 13.99 59.7 38.7
+ PointGST (ours) - 13.71 59.5 39.7
PointGPT-L [13] (baseline) NeurIPS 23 317.72 60.9 42.2
+ Linear probing - 15.48 59.9 39.1
+ IDPT [16] ICCV 23 24.93 60.0 41.7
+ Point-PEFT [18] AAAI 24 17.71 60.2 41.1
+ DAPT [17] CVPR 24 18.88 60.7 40.3
+ PointGST (ours) - 17.31 61.2 42.2
6.3 Part Segmentation
Part segmentation presents challenges in accurately predicting detailed class labels for each point. We apply the Point-BERT [9], Point-MAE [8], ACT [11], ReCon [37], and PointGPT-L [13] as baselines on the ShapeNetPart dataset. We incorporate the proposed point cloud spectral adapter (PCSA) into every layer of the models. The quantitative results are shown in Tab. V. Our method achieves highly competitive performance while utilizing significantly fewer trainable parameters than fully fine-tuning counterparts (i.e., baseline methods). Notably, the increase in trainable parameters is mainly due to the head part, but we still reduce overall trainable parameters and surpass other PEFT methods like IDPT [16], Point-PEFT [18] and DAPT [17].
Figure 6: t-SNE visualizations, along with trainable parameters (Params.) and overall accuracy (OA), are based on the ScanObjectNN PB_T50_RS dataset using a pre-trained PointGPT-L with different tuning strategies.
6.4 Scene-level Semantic Segmentation
We evaluate the proposed PointGST on the semantic segmentation task using the S3DIS dataset [81], as shown in Tab. VI. Our empirical observations suggest that previous PEFT methods [16, 18, 17] struggle with this task, likely due to the models being pre-trained on object-level datasets [93]. This results in sub-optimal performance when fine-tuned on a scene-level dataset. In contrast, our PointGST effectively utilizes intrinsic information during fine-tuning, bridging the performance gap between fully and efficiently fine-tuning strategies. For example, using ACT [11] as the baseline, PointGST achieves 67.6% mAcc and 57.4% mIoU with only 5.55M trainable parameters, substantially outperforming IDPT [16], Point-PEFT [18] and DAPT [17] by 3.5%/5.3%, 1.6%/2.8%, and 2.9%/2.9% in mAcc/mIoU, respectively. Similarly, when applied to Point-BERT [9], Point-MAE [8], ReCon [37], and PointGPT-L [13], PointGST consistently surpasses IDPT, Point-PEFT, and DAPT.
6.5 Scene-level Object Detection
Following the representative self-supervised methods [38, 9, 8, 11], we conduct object detection on the ScanNetV2 dataset [83]. Specifically, we replace the default encoder of a standard detector 3DETR [94] with the pre-trained backbone, and the results are presented in Tab. VII. The proposed PointGST outperforms all previous point cloud PEFT methods [16, 18, 17] under different baselines while with less trainable parameters. For instance, when applying ACT [11] as the baseline, our PointGST significantly surpasses the IDPT [16], Point-PEFT [18], and DAPT [17] by 4.1%, 4.3%, and 2.4% under the strict metric ($AP_{50}$) while using less trainable parameters. A similar phenomenon is also presented in the other baselines. These impressive results confirm the generalization ability of the proposed method on the scene detection task.
6.6 Point Cloud Completion
[…]
AdaPoinTr [95] (baseline) TPAMI 23 17.1 6.45 0.844
+ Linear probing - 0.4 6.87 0.822
+ IDPT [16] ICCV 23 1.7 6.75 0.831
+ Point-PEFT [18] AAAI 24 0.6 6.74 0.828
+ DAPT [17] CVPR 24 0.9 6.65 0.836
+ PointGST (ours) - 0.6 6.64 0.836